Home‎ > ‎Crew 136‎ > ‎

301-GreenHabReport

GreenHab Report
03/01/2014
Prepared by Benoit Pairet

The heater has been running: from 6pm to 9.30am

The pump is operating normally and the hydroponic tank water is full

This morning, I transplanted the beans and it went fine. I attached some
pictures for you to see.

I refilled the hydroponic tank along with Yoshini and naomi. I showed
them around the GreenHab and they are ready to take over.

I changed the pipe system as Chuck described to you this morning. It
seems to be working fine but it should be followed carefully to make
sure it does not do anything wrong.

The bucket under the hydro tank was empty and no transfer today

Greenhab readings for 5PM today are as follows:
Current Min Max
Temperature 66 F 50 F 84 F
Humidity 17 % 17 % 19 %

Greenhab readings at 9am today were as follows:

The pump did not work during the night since the reservoir was empty